OpenStudy (anonymous):

given the points A(1,-14) and B(9,2). Find the equation of the prepindicular bisector through the line segment AB

OpenStudy (dumbcow):

first find midpoint of AB ((1+9)/2, (-14+2)/2) = (5,-6) this is the point where the perpendicular line will go through Next find slope of AB slope = (2-(-14))/(9-1) = 16/8 = 2 slope of perpendicular line is opposite reciprocal of slope =-1/2 To get equation we need y-intercept use x,y points from midpoint above (5,-6) y=mx+b -6 = (-1/2)*5 + b b = -6 +5/2 = -7/2 y = -1/2 x -7/2